# TD-Orch: Scalable Load-Balancing for Distributed Systems with Applications to Graph Processing

### v1

@@
Abstract.
In this paper, we highlight a task-data orchestration abstraction that supports a range of distributed applications, including graph processing and key-value stores. Given a batch of tasks each requesting one or more data items, where both tasks and data are distributed across multiple machines, each task must get co-located with its target data (by moving tasks and/or data) and executed. We present TD-Orch, an efficient and scalable orchestration framework featuring a simple application developer interface. TD-Orch employs a distributed push-pull technique, leveraging the bidirectional flow of both tasks and data to achieve scalable load balance across machines even under highly skewed data requests (data hot spots), with minimal communication overhead. Experimental results show that TD-Orch achieves up to $2.7\times$ speedup over existing distributed scheduling baselines.
@@ Abstract.
graph processing; system orchestration; load balance; distributed-memory algorithms
conference: ACM SIGPLAN Annual Symposium on Principles and Practice of Parallel Programming; January 31–February 04, 2026; Sydney, Australiajournalyear: 2026copyright: none
1. Introduction
Distributed-memory systems, including multi-socket NUMAs (lameter2013numa), supercomputers (schneider2022exascale), compute-equipped disaggregated memory (intelagilex; calciu2021rethinking), processing-in-memory (upmem; mutlu2023primer), computational storage (samsungsmartssd; gu2016biscuit) and datacenters (soni2014novel; wu2015cloud; kumar2021resource) provide tremendous aggregated computation power and memory capacity. A broad class of distributed-memory applications, including key-value stores and graph processing systems, can be abstracted into repeated task–data orchestration problems. Here, a computational task maintains some context and attempts to access one or multiple data items located somewhere in the system. The goal is to co-locate the task context with (copies of) all target data on the same machine, execute a user-defined function $f$, and optionally write the result(s) back to the original data location(s).
Ideally, application developers only need to provide the high-level functions for $f$, getget, and $wb$ in Fig. 1 for their application, and the distributed system runtime ensures that tasks are executed correctly and efficiently on their target data. In this paper, we consider batch parallel task execution, where concurrent tasks are collected into batches that are executed in parallel, targeting high throughput. Thus, as shown in Fig. 1, the input to the orchestration interface (API) is an array of task contexts in a batch, where each context includes the address(es) of the task’s target data.
A simple example is provided in Algorithm 1, where conference attendance fees are charged from attendee’s deposit accounts residing in a key-value store. In this scenario, the application developer uses the API to specify the end-to-end outcome (i.e., the fees are deducted from the proper accounts), while relying on the task-data orchestration provided by the underlying system to achieve high throughput.
@@ 1. Introduction
    n: UINT, // Number of tasks in the batch
    tasks: array, // The array of task contexts
    f: function, // Execute using f(task, data)
    get: function, // get(pointer) -> data
    wb: function, // Writeback wb(pointer, data)
    merge_value: Merge-able operation in Sec 2.3
@@
Figure 1. Orchestration Function API. Task and available data start (roughly) evenly divided among the machines.
Input:
accounts: Distrib. data array of (conf_name, deposit) pairs
fees[$1:n$]: Distrib. task array of (data_addr, fee) pairs
1 Function f(task, data) :
2 return task.fee IF (data.conf_name == ”PPoPP”) ELSE 0
3Function get(data_addr) :
4 return accounts[data_addr]
5Function wb(data_addr, aggr_value) :
6 accounts[data_addr].deposit -= aggr_value
7Function ChargeConferenceFee(fees, accounts) :
8   Orchestration($n$, fees, $f$, getget, wbwb, ADD)
Algorithm 1 A Charge Conference Fee example
Providing such a task-data orchestration framework is challenging.
A key difficulty lies in achieving load balance in both communication and computation while keeping overhead low, thereby maximizing resource utilization and throughput. Each task requires all its data to be co-located on the same machine, by migrating the task and/or the data. Data targeted by many machines in a batch can become contention hot spots that violate load balance.
@@ 8   Orchestration($n$, fees, $f$, getget, wbwb, ADD)
In this paper, we propose TD-Orch, an orchestration framework that supports Fig. 1’s API, while being both theoretically grounded in scalable load balancing and practically efficient. Our key technique is distributed push-pull, which exploits the bidirectional movement of pushing tasks to data and pulling data to tasks—unlike most prior approaches that only push or pull. But judicious use of push and pull alone does not solve the hot spot problem, since neither pushing tasks to a machine holding hot data nor pulling the data to the tasks can avoid the high communication to/from that machine. Instead, TD-Orch detects hot data in each batch and employs novel (i) randomly-mapped reduction/broadcast communication trees and (ii) merge-able spillover meta-task structures. Together, these ensure (a) for low contention data items, requesting tasks’ contexts can be pushed directly to where the item resides (no tree hops), (b) for high contention data items, the set of machines needing the item are stored in a communication-bounded way, and (c) load balance and minimal communication overhead are achieved. Our evaluation of hash table workloads on 2–16 machines demonstrates a geomean speedup of $2.68\times$, $1.48\times$ and $2.70\times$ over the three baselines we study (§2.5), respectively.
To further demonstrate the power of TD-Orch, we apply TD-Orch to build a distributed graph processing system. Early distributed frameworks such as MapReduce (dean2008mapreduce) and the Pregel series (malewicz2010pregel; bu2014pregelix; han2015giraph) inspired many successors, including BLAS (bulucc2011combinatorial), Distributed GraphLab (low2012distributed), GraphMat (sundaram2015graphmat), Graphulo (gadepally2015graphulo), GraphX (gonzalez2014graphx), Mizan (khayyat2013mizan), Pegasus (kang2009pegasus), PowerGraph (gonzalez2012powergraph), PowerSwitch (xie2015sync), and Presto (venkataraman2013presto). We classify more recent frameworks developed in the past decade into two major families: (i) Graph-theoretic approaches: GeminiGraph (zhu2016gemini), Gluon (dathathri2018gluon), PowerLyra (chen2019powerlyra), FaasGraph (liu2024faasgraph), Phoenix (dathathri2019phoenix), ShenTu (lin2018shentu), and TopoX (li2019topox); and (ii) Linear-algebra-based designs: Graphite (mofrad2020graphite), GraphPad (anderson2016graphpad), GraphScope (fan2021graphscope), GraphTap (mofrad2019efficient), and LA3 (ahmad2018la3).
@@
We introduce TDO-GP, a distributed graph processing system atop TD-Orch. TDO-GP leverages TD-Orch as the primary mechanism for orchestrating task–data flow and employs implementation techniques to fully exploit the task–data layouts provided by TD-Orch. Like TD-Orch, TDO-GP is theoretically grounded, offering work-efficient bounds for the five graph algorithms implemented in this study. In addition to directly leveraging TD-Orch, we further optimize global communication, local computation, and the coordination between system components, tailoring these improvements for graph workloads to fully exploit the execution layout provided by TD-Orch. As a result, TDO-GP achieves geomean speedups of $4.1\times$ over the best prior open-source distributed graph processing systems designed for general graph problems (zhu2016gemini; mofrad2020graphite; ahmad2018la3) and up to $15\times$ in specific datasets.
In summary, the main contributions of this paper are:
  • We highlight task–data orchestration as an important practical problem in distributed applications, advocating a particular interface (Fig. 1).
@@
  • We design TD-Orch, a theoretically grounded and practically fast orchestration framework that achieves scalable load balance with low communication overhead.
@@
  • We develop TDO-GP, a distributed graph processing system built on TD-Orch and enhanced with tailored implementation techniques, which significantly outperforms prior systems (GeminiGraph, Graphite, LA3).
2. The Task-Data Orchestration Problem
In this section, we introduce an orchestration framework, TD-Orch, which aligns with the API in Figure 1. It yields optimal computation and communication costs while achieving scalable load balance. We will show that our design outperforms three prior methods for distributed load-balancing both in theory and in practice.
2.1. Problem Definition
In our abstract task-data orchestration problem for distributed-memory systems, there are $P$ machines, a batch of $n$ tasks to be executed on these machines, and data to be accessed by the tasks. The tasks are executed in stages and each task has a local context holder to store intermediate results. Data are partitioned into constant-sized data chunks. At the beginning of each stage, the tasks and the data chunks are evenly divided among the machines.
Each task requires accessing $O(1)$ data chunks (could be zero) that are stored outside its local context, and the task can be completed only after all the required data chunks are accessed. Each data chunk may be required by 0 to $n$ tasks.
@@ 2.1. Problem Definition
We allow for both tasks and data chunks to be moved between machines, and even routed through intermediate machines. Each movement of a task or a data chunk between two machines incurs a communication cost charged to both the sending and receiving machines. Executing a task incurs a computation work cost charged to the machine it executes on. Our goal is to orchestrate the tasks and data, i.e., to find an online computation scheduling and communication scheme that enables the stage to complete, such that the communication and computation work costs are (i) low and (ii) load-balanced across the machines.
@@
Load Balance. A key challenge in achieving load balance across the machines is the contention for the data chunks, since up to $n$ tasks may require the same data chunk. We seek task orchestration that is highly effective in practice and also with formal guarantees on its effectiveness regardless of the degree of data chunk contention. For the latter, we define the following goal:
@@
2.2. Prior Scheduling Strategies
Direct Pull. A naïve scheduling strategy would first eliminate duplicate requests for data chunks within each machine and then fetch all required data chunks to the corresponding tasks. This approach, commonly used in single-machine von Neumann architectures, is effective when the contention on each data chunk is low. However, data chunk contention can be high (e.g., in graph processing, where the contention may correspond to the degree of a vertex), resulting in high communication cost for machines holding “hot” data chunks and hence severe communication load imbalance.

### v2

@@
Abstract
In this paper, we introduce a task-data orchestration abstraction that supports a range of distributed applications, including graph processing and key-value stores. Given a batch of lambda tasks each requesting one or more data items, where both tasks and data are distributed across multiple machines, each task must be co-located with its target data (by moving tasks and/or data) and then executed. We present TD-Orch, an efficient and scalable orchestration framework featuring a simple application developer interface. TD-Orch employs a distributed push-pull technique, leveraging the bidirectional flow of both tasks and data to achieve scalable load balance across machines even under highly skewed data requests (data hot spots), with minimal communication overhead. Experimental results show that TD-Orch achieves up to $2.8\times$ speedup over existing distributed scheduling baselines.
@@ Abstract
1 Introduction
Distributed-memory computer systems, including supercomputers [schneider2022exascale], multi-socket NUMAs [lameter2013numa], compute-equipped disaggregated memory [intelagilex, calciu2021rethinking], processing-in-memory [upmem, mutlu2023primer], computational storage [samsungsmartssd, gu2016biscuit], and datacenters [soni2014novel, wu2015cloud, kumar2021resource] provide tremendous aggregated compute power and memory capacity. Motivated by the promise of these platforms, there has been a huge body of work over the past two decades on building high-performance distributed systems for applications such as graph processing [zhu2016gemini, chen2019powerlyra, gonzalez2014graphx, gonzalez2012powergraph, mofrad2020graphite], machine learning [moritz2018ray, narayanan2019pipedream, thorpe2021dorylus, abadi2016tensorflow, chen2015mxnet], key-value store [decandia2007dynamo, escriva2012hyperdex, chang2008bigtable, cooper2008pnuts], database [corbett2013spanner, shute2013f1, dageville2016snowflake, kim2025nocap], among many other application areas.
However, writing correct and high-performance applications on distributed-memory systems is notoriously challenging. There are mainly three strategies to implement distributed applications: (i) Use high-level interfaces such as Spark [zaharia2016apache], which are easy to program but leave notable performance on the table, since they are not optimized for specific end-to-end applications. (ii) Write low-level code using interfaces like the message passing interface (MPI), which introduces significant complexity in programming and system management. (iii) Use application-specific frameworks such as PowerGraph [gonzalez2012powergraph], which are optimized for certain domains but do not support new applications outside their scope. Consequently, these approaches often lead to wasted effort and repeated reinvention of techniques already embedded in other systems. For example, many graph frameworks repeatedly reimplement the ghost/mirror node design [gonzalez2012powergraph, zhu2016gemini, chen2019powerlyra, dathathri2019phoenix].
Ideally, one could leverage a high-level programming abstraction that enables users to write correct and high-performance distributed-memory applications while avoiding the complexity typically associated with writing high-performance distributed systems by hand, or the limitations of domain specific interfaces. For example, users could specify a description of the computations performed by the application while separating data from computational logic, e.g., by specifying: (i) which data to read from the system, (ii) the computation to perform on the retrieved values, and (iii) which results to write back. Such an abstraction crucially relies on an underlying system framework that automatically and efficiently executes the described application on the distributed system, handling all the underlying orchestration of computation and communication.
In particular, a common challenge of such orchestration is how to achieve load balance in both computation and inter-machine communication while minimizing overhead. Prior work tackles this in an ad-hoc fashion, e.g., through ideas like the above-mentioned mirror/ghost nodes in graph processing [zhu2016gemini, gonzalez2012powergraph], virtual servers in key-value stores [stoica2003chord, decandia2007dynamo], and workload-aware partitioning in databases [curino2010schism, quamar2013sword]. Despite differing implementations, the high-level ideas among these techniques are similar, suggesting that a clean abstraction could unify these approaches under a shared underlying system.
In this work, we build upon prior work on *-centric (e.g., node-centric, edge-centric, data-centric) abstractions by introducing a Task-Data Orchestration abstraction based on microservices-inspired lambda-tasks. Our lambda-task-centric abstraction can efficiently support a wide range of distributed applications, including graph processing, key-value stores, databases, failure-tolerant storage and others mentioned in §2.1. In our abstraction, users or application developers only need to specify the basic operation(s) to be performed by each lambda-task. The underlying system automates efficient orchestration of computation and communication, even under workload skew.
Abstraction. Our task-data orchestration abstracts any application that fits the following pattern. A batch of lambda-tasks (referred to as simply tasks in the rest of the paper) runs in parallel, where each task resembles a C++ closure, containing local metadata (called the task context) and a local lambda function representing the required computation, along with the locations of the inputs and outputs of the closure. The application proceeds in repeated bulk-synchronous parallel (BSP) [valiant1990bridging, valiant1990general, valiant1988optimally] orchestration stages, and in each stage each task (i) reads input data from outside its context, (ii) performs a local lambda computation using its context and the read values, and (iii) writes outputs to external locations.
As a concrete example, consider a distributed hash table that serves batches of search and update operations. Here, the tasks correspond to operations and the data reside in the hash-table storage. Each batch can be executed using a single orchestration stage that: (i) reads values from the hash table using task-specified keys, (ii) processes fetched values, and (iii) writes results back to the hash table.
Another concrete example is breadth-first search (BFS) for graphs. The algorithm proceeds in synchronous stages, maintaining a frontier of vertices $i$ steps away from the source during the $i$-th stage. Here, the tasks correspond to edges, and the data correspond to the distance values of vertices. During an orchestration stage, each edge (task) reads the distance values of its endpoints (data), checks whether one endpoint lies in the current frontier, and, if so, updates the distance of the other endpoint. This also fits into the orchestration abstraction of read–execute–write. A formal example of a BFS using this interface is shown in Algorithm˜1.
@@ 1 Introduction
tasks: An array of Task structures
wb: function, // wb(pointers[], values[])
struct Task {
pointer InputPointers[];
pointer OutputPointers[];
LocalContexts[]; // Per-Task local metadata
Lambda f; // f(InValues[]) -> OutValues[]
@@
Figure 1: The Task-Data Orchestration interface. Task and data are distributed among the same set of machines.
More formally, a single orchestration stage follows the interface in Fig. 1. Each task resembles a C++ closure: it contains some LocalContexts, a set of InputPointers indicating which external data items it will read, a local lambda function $f$ describing the per-task computation, and a set of OutputPointers indicating where returned results will be written. The lambda $f$ takes as input the local contexts and the values read from the InputPointers, and produces return values that will be written back to OutputPointers.
@@
System Design. In this paper, we propose TD-Orch, a new system that efficiently supports Fig. 1’s API, while being both theoretically grounded in scalable load balancing and practically efficient. Our key technique is distributed push-pull, which exploits the bidirectional movement of pushing tasks to data and pulling data to tasks—unlike most prior approaches that only push or pull. Judicious use of push and pull alone does not solve the hot spot problem, since neither pushing tasks to a machine holding hot data nor pulling the data to the tasks can avoid the high communication to/from that machine. Instead, TD-Orch detects hot data in each batch and employs novel (i) transit machines based on randomly-mapped communication trees and (ii) merge-able meta-task structures. Together, these ensure (a) for low contention data items, requesting tasks’ contexts can be pushed directly to where the item resides (no hops on a communication tree), (b) for high contention data items, the data will be sent to the tasks in a communication-efficient way, and (c) load balance and minimal communication overhead are achieved. Our evaluation of hash tables on 2–16 machines demonstrates a geomean speedup of $2.09\times$, $1.42\times$ and $2.83\times$ over three baselines we study (§4), respectively.
Combining Theory and Practice. One key contribution of this work is bridging theoretical analysis and empirical designs to achieve both load balance and low communication. Prior theory-driven approaches (e.g., [im2023massively, goodrich2011sorting]) achieve optimal bounds but adopt impractical techniques which lead to poor real-world performance, while empirical designs (e.g., [vora2014aspire, li2020pegasus, gog2016firmament, murray2013naiad]) lack theoretical guidance (no theoretical bounds in computation/communication costs), limiting their generalizability and robustness against skewed (or even adversarial) workloads. Our work fully redesigns theoretically optimal algorithms using practical implementation techniques such as distributed push-pull, achieving strong performance in both theory and practice. See §2.3 and §3.6 for further details of the comparison against prior work.
@@
Case Study: Distributed Graph Processing. Among the applications listed in §2.1, graph processing is likely one of the most important. To further demonstrate the power of TD-Orch, we develop TDO-GP, a distributed graph processing system atop TD-Orch. TDO-GP leverages TD-Orch as the primary mechanism for orchestrating task-data flow and employs implementation techniques to fully exploit the task-data layouts provided by TD-Orch. Like TD-Orch, TDO-GP is theoretically grounded, offering work-efficient bounds for the five graph algorithms implemented in this study. In addition to directly leveraging TD-Orch, we further optimize global communication, local computation, and the coordination between system components, tailoring these improvements for graph workloads to fully exploit the execution layout provided by TD-Orch. As a result, TDO-GP achieves geomean speedups of $4.1\times$ over the best prior open-source distributed graph processing systems designed for general graph problems [zhu2016gemini, mofrad2020graphite, ahmad2018la3, yan2015effective] and up to $15\times$ in specific datasets.
Overall, TDO-GP outperforms two major families of prior graph systems: compared with graph-algorithm-based systems (e.g., [zhu2016gemini, chen2019powerlyra, yan2015effective]), it provides significantly better provable and empirical load balance through the use of transit machines; and compared with linear-algebra-based approaches (e.g., [mofrad2020graphite, ahmad2018la3, anderson2016graphpad]), it achieves stronger graph-theoretic guarantees and higher performance by better exploiting graph sparsity. Moreover, the interface enables user implementation of a complex graph algorithm in just 70 lines of C++ code, versus prior 400 lines in [zhu2016gemini] (see §5).
Main Contributions. In summary, in this paper:
  1. (1)
    We are the first to cleanly define task-data orchestration, a fundamental execution pattern in distributed applications, and to advocate a concrete interface for it (Fig. 1).
  2. (2)
@@
    We design TD-Orch, a theoretically grounded and practically fast orchestration system that achieves scalable load balance with low communication overhead.
  3. (3)
@@
    We develop TDO-GP, a distributed graph processing system built on TD-Orch and enhanced with tailored implementation techniques, which significantly outperforms prior state-of-the-art systems [zhu2016gemini, mofrad2020graphite, ahmad2018la3, yan2015effective].
2 Preliminaries
2.1 TD-Orch: Supported Applications
Many distributed applications can be abstracted using the task-data orchestration interface. We show here several examples.
Graph Processing. Many vertex- and edge-centric graph algorithms share a common execution pattern: a vertex array stores intermediate values; in each stage, a subset of edges are activated, read the values of their endpoints, perform a computation using these vertex values and the edge weight, and write the result back to the vertex array. Examples include breadth-first search (BFS), shortest paths, betweenness centrality, connectivity, PageRank, and spanning forests.
All such graph applications can be expressed within the task-data orchestration framework. Example pseudocode [^1] This example demonstrates how graph algorithms can be written using the orchestration interface. It is intentionally simple but not work-efficient. Work-efficient implementations rely on auxiliary structures such as DistVertexSubset, introduced later in §5. for BFS is shown in Algorithm˜1, where Tasks correspond to edges and Data correspond to vertex values—the distance from the source node in BFS. At each stage, lambda $f$ is applied to each edge. It checks whether the source vertex of the edge was activated in the previous round, having a distance of $\textsc{round}-1$. If so, it writes distance round to the destination vertex, provided it has not been written already.
Input: $G=(V,E)$, where $|V|=n$, $|E|=m$; StartVertex
Output: Dist[$1:n$]: Distrib. vertex value array
[…]
2 VertexUpdated = 1
3 round = 1
4 Struct Task
5 LocalContext: Edge (u, v)
6  InputPointer: pointer(dist[u])
7  OutputPointer: pointer(dist[v])
8  Function f(value_u: int) :
9     return round If (value_u == round - 1) Else -1
10Function WriteBack(dist_ptr: pointer, value_v: int) :
11  If *dist_ptr == -1 and value_v != -1 then
12     *dist_ptr = value_v
13     VertexUpdated = 1
14Function BFS($G$, StartVertex: int) :
15  Dist[StartVertex] = 0
16  Initialize Task array tasks[1:m] using G.E[1:m]
17  While VertexUpdated > 0 do
18     VertexUpdated = 0
19     Orchestration(tasks, WriteBack)
20     round += 11
21 return Dist[$1:n$]
Algorithm 1 Breadth-First Search (BFS)
Key-Value Store. Many key-value store abstractions naturally fit within the task-data orchestration model. A straightforward example is a (distributed) hash table, where reading and updating a batch of items can be expressed as a one-stage orchestration by defining $f$ as the per-item operation. Further details are provided in §6.
Similarly, many ordered indexes (e.g., B-trees, radix trees) can be represented within this abstraction by decomposing each search/update query into multiple orchestration stages. In each stage, a single level of the index/tree is traversed: the corresponding internal node is retrieved, and $f$ performs the search within the node to identify the appropriate child. This process continues until the leaf is reached.
Transactional/Analytical Databases. Transactional (OLTP) and analytical (OLAP) databases share structural similarities with key-value stores, often relying on tree-based index structures. Their index traversals can be orchestrated in the same way as key-value lookups to support search and range scan queries. In addition, OLTP/OLAP systems maintain version chains, which resemble concurrent bucket-based hash tables and are likewise supported by the orchestration framework.
Vector Search. High-dimensional approximate nearest neighbor (ANN) search is a key component in vector search. Many state-of-the-art implementations use bucket-based designs [tian2023db], graph-based approaches [jayaram2019diskann, malkov2018efficient], or hybrid schemes [munoz2019hierarchical, manohar2024parlayann] that maintain a nearest-neighbor graph at higher levels to guide searches toward candidate buckets for detailed vector comparisons. The bucket traversal naturally fits the orchestration framework, similar to a key-value store, while the graph traversal resembles a vertex-centric graph algorithm, which can also be expressed using orchestration.
Fault-Tolerant Storage. Our orchestration framework can also support RAID [chen1994raid] or erasure coding [balaji2018erasure] in fault-tolerant storage. By setting InputPointers to fetch the original data, WB to write back to the target location at OutputPointers, and $f$ to perform the RAID or erasure-coding computation, a redundancy-based erasure-resilient storage abstraction can be efficiently implemented across the machines.
2.2 System Modeling
The task-data orchestration interface maps the execution flow onto a distributed system whose details—such as scheduling and physical data partitioning—are not exposed to the user.
Model. To provide better quantified analysis on our and prior designs, we use the Bulk-Synchronous Parallel (BSP) model [valiant1990bridging, valiant1990general, valiant1988optimally]. The modeling assumes the system consists of $P$ machines, each with a processing unit and a local memory without shared global memory. All machines are connected by a point-to-point communication network. The system operates in rounds (or supersteps), where periodic barrier synchronizations are required, similar to MPI.
BSP accounts for computation time (maximum work) and communication time (maximum data exchanged) across all machines. Since the metrics use “maximum”, maintaining load balance is crucial. Readers can refer to Appendix A for more details of the model.
Data Storage. Data are partitioned into data chunks of granularity $B$ words, analogous to cache-line or page storage to exploit locality. Each chunk is placed on a random machine, providing adversary-resistant load balance in both storage and access patterns [sanders1996competitive]. Such randomized placement is commonly used to ensure robustness in high performance [wang2023provably, kang2022pimtreevldb, shan2023explore, luo2023smart].
Executed Tasks. A batch of $n$ tasks, each with a context size of at most $\sigma$ words, will be executed during an application. At initialization, the tasks are evenly distributed across the machines. During each orchestration stage, each task requires access to at most $D$ data chunks (possibly zero) stored outside its local context, and a task completes only after all required data chunks have been fetched, processed and written back.
@@ 2.2 System Modeling
Design Goals. Both tasks and data chunks are allowed to be moved between machines, and even routed through intermediate machines. Our goal is to orchestrate tasks and data—i.e., to design an online scheduling scheme that (i) minimizes the amounts of communication and computation, and (ii) ensures load balance across machines to maximize throughput.
@@
Load Balance. A key challenge in achieving load balance across the machines is the contention for the data chunks, since up to $n$ tasks may require the same data chunk. We seek orchestration that is highly effective in practice and also with formal guarantees regardless of the degree of data chunk contention. For the latter, we define the following goal:
@@
2.3 Scheduling Strategies
As described in §2.2, task-data orchestration basically seeks an execution scheme in which $n$ task contexts, each of size $O(\sigma)$, access $O(nD)$ data chunks, each of size $B$. The ideal communication time of $O(nD\min\{B,\sigma\}/P)$ is unachievable in practice using only online scheduling algorithms. Instead, prior work typically adopts one of the following strategies.
Direct Pull. A naïve scheduling strategy first eliminates duplicate requests for data chunks within each machine, then fetches all required chunks to the corresponding tasks. This approach, common in remote direct memory accesses (RDMA), works well when contention (reference count) on each chunk is low. However, contention can be high—for example, in graph processing, contention can be caused by extremely high-degree vertices—leading to high communication on machines storing “hot” chunks and severe load imbalance. In the worst case, all $n$ tasks may request the same chunk, incurring $O(DPB/\min\{D,P\})$ communication time and competitive ratio of $O\left(\frac{P^{2}B}{n\min\{B,\sigma\}\min\{D,P\}}\right)$, which is prohibitive as $P$ grows.